OpenStudy (anonymous):

In this question you will derive a general formula for the distance of a point from a line. Let P be the point (p,q) and L the line y=mx+b. The slope of L is m The slope of a line perpendicular to L is -1/m The line through P perpendicular to L can be written as y=sx+c where s is: -1/m and c is: q+(p/m) That line intersects L in the point Q=(u,v), where u is: __________ and v is: ____________ The distance of P and Q is ________

ganeshie8 (ganeshie8):

solve below equations for x and y : y = mx + b y = sx + c

ganeshie8 (ganeshie8):

can you ?

OpenStudy (anonymous):

y = mx+sx+b+c ? @ganeshie8

OpenStudy (anonymous):

over 2 sorry so y = (mx+sx+b+c )/2